OpenStudy (narissa):

y=6 x=20 what is the constant of variation in decimal form

OpenStudy (narissa):

@texas cic 101

OpenStudy (narissa):

@texaschic101

OpenStudy (narissa):

its .3 or 1.8?

OpenStudy (amorfide):

assuming that y varies directly with x then the regular equation for the constant of variation will be y=kx where k is a constant to figure this out we substitute out values y=6 and x=20 to get 6=20k 6/20=k k=3/10 k=0.3 constant of variation is 0.3

OpenStudy (narissa):

yeah I got it half way right!

OpenStudy (narissa):

this one says y=6 and x=-36 what is direct variation equation y=6x, y=1/6, y=6x or y=-1/6